# AdvLoRA: Adversarial Low-Rank Adaptation of Vision-Language Models

### v1

@@ 4.3. Performance Comparisons
In this section, we conduct a comparative analysis between our proposed AdvLoRA and five baselines across two cross-modal retrieval tasks using four datasets. Specifically, we perform adversarial adaptation based on the PGD-3 (Madry et al., 2018) attack to all methods and then evaluate their performance under the condition of adversarial attack data and natural data.
Firstly, for image-text retrieval, we conducted experiments on adversarial attacked data for both Flickr30K and MSCOCO, as shown in Table 4 and Table 3. From these experiments, we draw two important conclusions as follows.
1) After adversarial adaptation, AdvLoRA outperforms all other baselines when faced with adversarial attacks. Notably, on MSCOCO, AdvLoRA surpasses all other PEFT methods by 12.17% and exceeds FFT by 2.47%, while using approximately 100x fewer tunable parameters than FFT.
@@
2) AdvLoRA demonstrates enhanced adversarial robustness on larger datasets, highlighting the significant potential of PEFT methods in improving model robustness against adversarial attacks. Specifically, on the relatively smaller dataset Flickr30K, the performance of various baselines after adversarial adaptation is comparable and does not show a significant increase in robustness. However, on the larger dataset MSCOCO, FFT achieves considerable adversarial robustness, yet it still lags behind AdvLoRA. These results benefit not only from the design of AdvLoRA in terms of clustering reparameterization and parameter alignment but also indicate that the effectiveness of adversarial adaptation improves with the increase of adaptation data.
Secondly, for video-text retrieval, we conducted experiments on adversarial attacked data for both Didemo and MSR-VTT datasets, as shown in Table 6 and Table 5. From these experiments, we draw two conclusions from the image-text retrieval as follows.
1) AdvLoRA achieves excellent adversarial robustness on video data, surpassing all other baselines. In DiDeMo, AdvLoRA slightly outperforms Uniadapter while using 7x fewer parameters. On MSR-VTT, AdvLoRA enhances the model’s adversarial robustness by 39.16% and significantly exceeds the other baselines.
2) AdvLoRA demonstrates better adversarial robustness on larger datasets. Specifically, on the relatively smaller dataset DiDeMo, the performance of various baselines after adversarial adaptation is comparable, and the robustness improvement is not significant. However, on the larger dataset MSR-VTT, the Uniadapter method achieves considerable adversarial robustness but is still inferior to AdvLoRA, and it uses 7x many parameters. Such results are attributed to the design of AdvLoRA in terms of clustering reparameterization and parameter alignment. It indicates that the effectiveness of adversarial adaptation improves with the increase of adaptation data.
@@
Thirdly, we conducted experiments on natural data from four datasets, and Table 7 presents the results for MSCOCO. The complete results on other datasets are provided in the Appendix. From these experiments, we draw a significant conclusion as follows.
Adversarial adaptation can degrade the performance of the model on natural data. For instance, a comparison between Table 7 and Table 1 reveals that, except LP and LoRA, all other models experience a decline in performance after adversarial adaptation. However, the AdvLoRA method still achieves competitive results on MSCOCO. This can be attributed to AdvLoRA’s ability to learn semantically invariant feature representations. The reason for the lack of performance degradation in LP and LoRA may be due to their low sensitivity to adversarial adaptation, leading to an ineffective adaptation process. As shown in Table 3, LP and LoRA do not acquire improved adversarial robustness after adversarial adaptation.
Table 7. Natural experiment with adversarial adaptation on MSCOCO. “Nat” donates natural images. An asterisk (*) indicates that adversarial adaptation has been performed.
Method Tunable Para. MSCOCO TR MSCOCO IR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT*+Nat 223M 57.28 78.92 86.36 74.17 48.52 75.04 83.77 69.11 71.65
BLIP+LoRA*+Nat 2.8M 70.76 90.44 94.68 85.29 56.39 80.38 87.48 74.75 80.02
BLIP+LP*+Nat 0.5M 72.58 91.2 95.22 86.33 57.15 80.93 88.05 75.38 80.86
UniAdapter*+Nat 19.5M 55.62 81.24 89.02 75.29 45.06 72.99 82.61 66.89 71.09
Aurora*+Nat 0.3M 70.92 89.39 93.94 84.74 54.38 79.38 86.88 73.54 79.15
AdvLoRA*+Nat 2.8M 70.58 90.42 94.54 85.18 56.36 80.35 87.29 74.67 79.92
4.4. Ablation Study
Figure 3. Ablation study. A, B, C, D denotes LoRA, LoRA with parameter clustering, LoRA with parameter clustering and alignment, and AdvLoRA, respectively.
In this section, we conduct an ablation study on AdvLoRA to demonstrate the effectiveness of the proposed clustering reparameterization, parameter alignment, and adaptive parameter update strategy on Didemo, and the results are presented in Figure 3. We draw the following conclusions.
The techniques proposed, including clustering reparameterization, parameter alignment, and adaptive parameter updates, are effective. Moreover, as evidenced by Figure 3, the model achieves optimal adversarial robustness when these methods are collectively employed.
4.5. Adaptation Efficiency and Storage Cost
In this section, we conduct an analysis and comparison of the adaptation efficiency and storage cost associated with AdvLoRA. Table 8 illustrates the relative training GPU hours and GPU memory cost, where the time (or memory) of FFT is taken as one unit. The following conclusions can be drawn. 1) In terms of time overhead, AdvLoRA does not exhibit a pronounced advantage, but it outperforms Aurora and FFT. It is noteworthy that the adaptation process of models based on online weight decomposition, such as Aurora, requires more time than FFT. In contrast, AdvLoRA has a smaller time overhead due to the completion of only one offline clustering reparameterization and parameter alignment before adaptation. 2) In terms of memory overhead, AdvLoRA surpasses Aurora and FFT. Aurora again experiences a higher memory cost than FFT due to its heavier online decomposition. 3) Overall, AdvLoRA, without any additional constraints on training time and memory, can be considered an excellent adversarial adaptation method to enhance the adversarial robustness of VLMs.
Table 8. Comparison on the training time and GPU memory.
Method Tuable Para. Time Mem.
BLIP+FFT 223M 1.00 1.00
BLIP+LoRA 2.8M 0.91 0.85
BLIP+LP 0.5M 0.79 0.67
Uniadapter 19.5M 0.93 0.77
Aurora 0.3M 1.05 1.04
AdvLoRA 2.8M 0.94 0.85
4.6. Hyperparameter Sensitivity Analysis
In this section, we conduct a sensitivity analysis on the rank size of AdvLoRA on Flickr30K. We set a series of values for the rank, namely 8, 10, 16, 32, and 64, and the results are presented in Figure 4 (a). AdvLoRA is not sensitive to the rank size, allowing us to select an appropriate rank according to our needs to reduce the cost of adaptation.
(a) Sensitivity Analysis
(b) Convergence Analysis
Figure 4. Analysis experiments of our proposed method
4.7. Loss Convergence Analysis
In this section, we conduct a convergence analysis experiment between AdvLoRA and LoRA on Flickr30K. The results are presented in Figure 4 (b). Analysis of the experimental results we can draw the following conclusion. 1) AdvLoRA demonstrates superior convergence over LoRA in the adversarial adaptation process, achieving a significantly reduced loss level. 2) AdvLoRA accelerates the convergence of adversarial adaptation more effectively than LoRA. These efficiencies and effectiveness can be attributed to the design of clustering reparameterization, parameter alignment, and adaptive parameter update strategy.
@@ 5. Conclusion
In this paper, we aim to alleviate the security risks in the Vision-Language Models (VLMs). First of all, we show the vulnerability of VLMs with various adaptation methods under adversarial attacks via extensive experiments. Besides, as the sizes of VLMs increase, simply applying the conventional adversarial adaptation methods to VLMs easily leads to 1) unpromising adversarial robustness and 2) tremendous parameter and training costs. From these motivations, a novel parameter-efficient adversarial adaptation method named AdvLoRA is proposed with parameter clustering, parameter alignment, and adaptive parameter update. Extensive experiments demonstrate the effectiveness and efficiency of AdvLoRA. This result reveals the intrinsic low-rank property that emerges during the adversarial adaptation process. Our proposed technique, which involves clustering reparameterization and parameter alignment, has been instrumental in facilitating the adaptation process. We have thereby offered a novel perspective for researchers in the field of security within the broader context of AGI.
However, in this paper, we merely use the simple PCD-3 for the adversarial attacks. In the future, it is worth adopting more challenging attacks and further optimizing the memory and computational budget during the adaptation process.
@@ Acknowledgements.
This work is supported by the Ministry of Science and Technology of China under Grant No, 2020AAA0108401, and the Natural Science Foundation of China under Grant Nos. 72225011 and 72293575. Note that Xiaolong Zheng is the corresponding author of this article.
Appendix A Datasets
Appendix B Baselines
  • BLIP-FFT is a conventional adaptation technique that enhances the performance of BLIP for specific downstream tasks by retraining and updating full parameters in downstream tasks.
  • BLIP-LP is an adaptation technique that involves adding and training a linear layer on top of the frozen pre-trained model BLIP to adapt to specific downstream tasks.
  • BLIP-LoRA is a Parameter-Efficient Fine-Tuning (PEFT) technology that adapts BLIP by introducing low-rank adapters to capture task-specific information, allowing for efficient adaptation to downstream tasks with minimal tunable parameter updates.
Appendix C Hyperparameter Setting
We present the hyperparameter setting in Table 9.
Table 9. Hyperparameter setting
config Image-text Retrieval Video-text Retrieval
Flickr30K MSCOCO Didemo MSR-VTT
optimizer AdamW AdamW AdamW AdamW
learning rate 1e-5 1e-5 1e-4 1e-4
schedule cosine decay cosine decay cosine decay cosine decay
training batchsize 16 16 8 8
inference batchsize 32 32 8 8
frames - - 16 16
attack ratio - - 20% 20%
epochs 5 5 5 5
training input 384 384 8*224 8*224
inference input 384 384 16*224 16*224
adversarial type PGD-3 PGD-3 PGD-3 PGD-3
attack alpha 1/255 1/255 1/255 1/255
PGD-epsilon 1/255 1/255 1/255 1/255
rank 10 10 10 10
adaptive weight 1e-3 1e-3 1e-3 1e-3
weight norm learning rate 1e-3 1e-3 1e-3 1e-3
Appendix D Vulnerability to Adversarial Attacks
In this section, we present the vulnerability results on Flickr30K and DiDeMo in Table 10 and Table 11. We have the conclusion that BLIP adapted by different methods are highly susceptible to adversarial perturbations, which is similar to that in the main text.
Table 10. Vulnerability experiment on Flickr30K. “FFT” and “LP” denoting full fine-tuning and linear probe. “Nat” and “Att” donate natural images and adversarially attacked images. “TR” and “IR” donate text-to-image retrieval and image-to-text retrieval.
Method Tunable Para. Flickr30K TR Flickr30K IR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT+Nat 223M 72.8 90.8 95.5 86.37 63.4 86.58 92 80.66 83.52
BLIP+FFT+Att 21.1 38.4 46 35.16 21.96 42.62 51.68 38.58 36.87(-46.65%)
BLIP+LP+Nat 0.5M 89 98.5 99.5 95.67 78.32 94.34 96.98 89.88 92.78
BLIP+LP+Att 55.9 76 81.7 71.2 49.3 70.82 77.48 65.87 68.54(-24.24%)
BLIP+LoRA+Nat 2.8M 87 98.1 99.5 94.87 72.9 93.9 96.84 87.88 91.38
BLIP+LoRA+Att 71.6 92.1 95.5 86.4 60.62 85.92 91.18 79.24 82.82(-8.56%)
UniAdapter+Nat 19.5M 96.7 99.7 100 98.8 86.18 97.34 98.82 94.11 96.46
UniAdapter+Att 70.2 85.5 89.5 81.73 61.26 80.26 86.3 75.94 78.84(-17.62%)
Aurora+Nat 0.3M 96.7 99.8 100 98.83 85.76 97.24 98.72 93.91 96.37
Aurora+Att 69.4 84.7 88.4 80.83 60.98 80.64 86.22 75.95 78.39(-17.98%)
Table 11. Vulnerability experiment on Didemo. “TR” and “VR” donate text-to-video retrieval and video-to-text retrieval, respectively.
Method Tunable Para. Didemo TR Didemo VR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT+Nat 223M 30.51 55.63 66.4 50.85 32.8 58.52 68.49 53.27 52.06
BLIP+FFT+Att 12.66 26.32 35.39 24.79 14.56 31.7 40.58 28.95 26.87(-25.19%)
BLIP+LP+Nat 0.5M 25.32 44.77 53.24 41.11 26.82 50.25 58.42 45.16 43.14
BLIP+LP+Att 23.13 45.86 53.54 40.84 26.02 47.06 57.03 43.37 42.11(-1.03%)
BLIP+LoRA+Nat 2.8M 36.79 63.21 72.28 57.43 34.1 62.41 73.08 56.53 56.98
BLIP+LoRA+Att 33.2 57.43 66.7 52.44 32.7 56.73 68.1 52.51 52.48(-4.51%)
UniAdapter+Nat 19.5M 32.8 60.02 71.19 54.67 9.97 28.32 40.38 26.22 40.45
UniAdapter+Att 27.02 52.14 64.01 47.72 9.27 24.83 36.69 23.6 35.66(-4.79%)
Aurora+Nat 0.3M 35.59 63.61 73.08 57.43 37.49 63.01 72.68 57.73 57.58
Aurora+Att 30.31 52.94 64.11 49.12 31.21 54.74 64.61 50.19 49.66(-7.92%)
Appendix E Performance on Natural data
In this section, we present the performance results of natural data on Flickr30K, DiDeMo and MSR-VTT in Table 12, Table 13 and Table 14. We have the following conclusion similar to that in the main text.
Adversarial adaptation can degrade the performance of the model on natural data. However, the AdvLoRA method still achieves competitive results on these datasets. This can be attributed to AdvLoRA’s ability to learn semantically invariant feature representations. The reason for the lack of performance degradation in LP and LoRA may be due to their low sensitivity to adversarial adaptation, leading to an ineffective adaptation process.
Table 12. Natural experiment with adversarial adaptation on Flickr30K. “Nat” donates natural images. An asterisk (*) indicates that adversarial adaptation has been performed. The best results are displayed in bold, while the second-best results are underlined.
Method Tunable Para. Flickr30K TR Flickr30K IR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT+Nat 223M 72.8 90.8 95.5 86.37 63.4 86.58 92 80.66 83.51
BLIP+LoRA+Nat 2.8M 96.9 99.9 100 98.93 86.72 97.78 98.82 94.44 96.69
BLIP+LB+Nat 0.5M 89 98.5 99.5 95.67 78.32 94.34 96.98 89.88 92.77
UniAdapter+Nat 19.5M 96.7 99.7 100 98.8 86.18 97.34 98.82 94.11 96.46
Aurora+Nat 0.3M 96.7 99.8 100 98.83 85.76 97.24 98.72 93.91 96.37
AdvLoRA+Nat 2.8M 96 99.7 100 98.57 85.68 97 98.64 93.77 96.17
Table 13. Natural experiment with adversarial adaptation on Didemo. “Nat” donates natural videoes. An asterisk (*) indicates that adversarial adaptation has been performed. The best results are displayed in bold, while the second-best results are underlined.
Method Tunable Para. Didemo TR Didemo VR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT+Nat 223M 30.51 55.63 66.4 50.85 32.8 58.52 68.49 53.27 52.06
BLIP+LoRA+Nat 2.8M 36.79 63.21 72.28 57.43 34.1 62.41 73.08 56.53 56.98
BLIP+LB+Nat 0.5M 25.32 44.77 53.24 41.11 26.82 50.25 58.42 45.16 43.14
UniAdapter+Nat 19.5M 32.8 60.02 71.19 54.67 9.97 28.32 40.38 26.22 40.45
Aurora+Nat 0.6M 35.59 63.61 73.08 57.43 37.49 63.01 72.68 57.73 57.58
AdvLoRA+Nat 2.8M 32.1 60.72 69.39 54.07 35.29 59.82 71.18 55.43 54.75
Table 14. Natural experiment with adversarial adaptation on MSR-VTT. “Nat” donates natural images. An asterisk (*) indicates that adversarial adaptation has been performed. The best results are displayed in bold, while the second-best results are underlined.
Method Tunable Para. MSR-VTT TR MSR-VTT VR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT+Nat 223M 20.3 41.3 53.8 38.47 23.4 48.4 60.8 44.2 41.33
BLIP+LoRA+Nat 2.8M 47.2 71.4 80.5 66.36 45.8 70.7 80.3 65.6 65.98
BLIP+LB+Nat 0.5M 40.3 63.2 72.0 58.5 41.8 63.7 71.6 59.03 58.77
UniAdapter+Nat 19.5M 42.4 68.4 77.4 62.73 42.9 68.4 78.3 63.2 62.97
Aurora+Nat 0.6M 45.1 69.7 79.4 64.73 44.2 68.5 77.8 63.5 64.12
AdvLoRA+Nat 2.8M 47.1 71.8 81.9 66.93 47.5 71.2 79.9 66.2 66.57
Appendix F Case Study
In this section, we conduct a case study on MSR-VTT, as illustrated in Figure 5. It can be observed that AdvLoRA achieves robust retrieval performance under adversarial attacks.
(a) Positive Sample Pairs
Figure 5. Case study of MSR-VTT. We sample and visualize eight frames from the videos. The frames with the devil denote that they are under the adversarial attacks. The first and second texts are the output of AdvLoRA and Aurora, respectively.

### v3

@@ 4.2. Performance Comparisons
We compare our method with five baselines across two cross-modal retrieval tasks using four datasets. Specifically, we perform adversarial adaptation based on the PGD-3 attack for all methods and evaluate their performance under adversarial conditions.
@@
Evaluation on Image-Text Retrieval. We conducted experiments on adversarially attacked data for both MSCOCO and Flickr30K, as presented in Tab.2 and Tab.3. The results lead us to two key conclusions: (1) Performance Under Adversarial Attacks. After adaptation, AdvLoRA outperforms all other baselines when confronted with adversarial attacks. On MSCOCO, AdvLoRA surpasses other PEFT methods by 12.17% and exceeds FFT by 2.47%, using approximately 100 times fewer tunable parameters. (2) Enhanced Robustness with Larger Datasets. AdvLoRA shows improved robustness on larger datasets, highlighting the potential of PEFT methods to bolster model resilience. On the smaller Flickr30K dataset, the performance of various baselines post-adaptation remains comparable, with no significant robustness increase. However, on MSCOCO, FFT achieves notable robustness but still trails behind AdvLoRA. These results emphasize AdvLoRA’s advantages in clustering reparameterization and parameter alignment.
Evaluation on Video-Text Retrieval. We conducted experiments on adversarially attacked data for both DiDeMo and MSR-VTT, as presented in Tab.6 and Tab.7. Our findings lead to the following conclusions: (1) Adversarial Robustness of AdvLoRA. AdvLoRA demonstrates excellent adversarial robustness on video data, surpassing all other baselines. In DiDeMo, AdvLoRA slightly outperforms UniAdapter while utilizing seven times fewer parameters. On MSR-VTT, AdvLoRA enhances the model’s adversarial robustness by 39.16%, significantly exceeding other baselines. (2) Improved Robustness with Larger Datasets. AdvLoRA shows better adversarial robustness on larger datasets. On the smaller DiDeMo dataset, the performance of various baselines after adversarial adaptation remains comparable, with minimal improvements in robustness. In contrast, on the larger MSR-VTT dataset, while UniAdapter achieves notable adversarial robustness, it still falls short of AdvLoRA’s performance, using seven times more parameters. These results can be attributed to AdvLoRA’s design, particularly in terms of clustering reparameterization and parameter alignment. These results further highlight the design advantages of AdvLoRA in clustering reparameterization and parameter alignment.
Table 4. Ablation study on AdvLoRA components. PC, PA, and PU denote parameter clustering, parameter alignment, and adaptive parameter update, respectively. The metrics TR@Mean and VR@Mean denote the mean recall for text-to-video and video-to-text tasks, respectively, while R@Mean represents the average of TR@Mean and VR@Mean.
Setting PC PA PU TR@Mean VR@Mean R@Mean
Baseline 54.37 54.07 54.22
a 55.36 56.29 55.83
b 57.85 57.21 57.53
c 57.69 58.42 58.01
4.3. Ablation Study
Effect of Proposed Module. To demonstrate the effectiveness of the proposed module, we conduct an ablation study on the components of AdvLoRA using the Didemo dataset and summarize our results in Tab. 4. Firstly, we explored the impact of the parameter clustering technique. Our experiments, comparing model variant (a) with the baseline, indicate a notable improvement when leveraging knowledge from pre-trained models through parameter clustering. Moreover, the performance gains observed in model variants (a), (b), and (c) highlight that incorporating parameter alignment and adaptive parameter update methods on top of parameter clustering significantly enhances the adversarial adaptability of VLMs. Finally, using all the proposed methods together, i.e. AdvLoRA, results in the best overall performance, achieving a mean recall of 58.01.
Sensitivity Analysis of Rank Size. To demonstrate the impact of different rank sizes on AdvLoRA, we trained a series of models with varying ranks on Flickr30K. The results are shown in Fig. 4. We find that as the rank size increases, the number of tunable parameters increases, but AdvLoRA’s performance remains stable, while computational cost rises. Therefore, during adversarial adaptation, performance is not sensitive to rank size, allowing us to choose a smaller rank without significant loss in performance.
Figure 3. Sensitivity Analysis.
Figure 4. Loss Analysis.
Loss Convergence Analysis. To explore the effectiveness of the proposed method in terms of loss convergence, we examine two model variants, LoRA and AdvLoRA, on the Flickr30K dataset. The results are presented in Fig. 4. The experimental findings indicate that (1) AdvLoRA achieves a lower loss early in the training process, demonstrating that techniques we proposed such as parameter clustering can serve as an effective ”warm-up,” helping VLMs establish a favorable starting point during adversarial adaptation; (2) as training progresses, AdvLoRA successfully navigates to a more effective gradient descent trajectory, leading to more efficient model training. In contrast, LoRA appears to get trapped in a local optimum. This highlights the superiority of our proposed method in facilitating better convergence of VLMs.
Table 5. Comparison on the training time and GPU memory.
Method Tuable Para. Time Mem.
BLIP+FFT 223M 1.00 1.00
BLIP+LoRA 2.8M 0.91 0.85
BLIP+LP 0.5M 0.79 0.67
Uniadapter 19.5M 0.93 0.77
Aurora 0.3M 1.05 1.04
AdvLoRA 2.8M 0.94 0.85
@@ 4.3. Ablation Study
Efficiency and Storage Cost Analysis. To investigate the cost-effectiveness of the proposed method, we analyze the relative training GPU hours and GPU memory costs of AdvLoRA compared to five baseline models on Flickr30K. The results are presented in Tab. 5, where the time (or memory) of FFT is normalized to one unit. The findings indicate that: (1) AdvLoRA exhibits relatively low parameter count and memory overhead; (2) AdvLoRA incurs longer time costs due to the necessity of offline clustering reparameterization and parameter alignment prior to adaptation.
Table 8. Additional attack types on the MSCOCO dataset. “TR” and “IR” donate text-to-image retrieval and image-to-text retrieval.
Method Attack TR@Mean IR@Mean Mean
LoRA PGD-3 61.01 54.41 57.71
AdvLoRA PGD-3 82.40 69.83 76.12
AdvLoRA PGD-20 81.65 69.13 75.39
AdvLoRA FGSM 84.44 72.21 78.32
AdvLoRA BIM 83.25 69.56 76.41
AdvLoRA SA 87.40 75.83 81.62
AdvLoRA ZOO 84.17 73.83 79.00
Performance on More Attacks. To demonstrate the robust generalization of the proposed method, we apply additional attacks to the AdvLoRA model, which has undergone adversarial adaptation on the MSCOCO dataset. These include white-box attack methods (FSGM (Goodfellow et al., 2014), PGD-20 (Madry, 2017), BIM (Kurakin et al., 2016)) and black-box methods (Zeroth Order Optimization (ZOO) (Chen et al., 2017), SquareAttack with 500 queries (SA) (Andriushchenko et al., 2020)), as shown in Tab. 8. Note that AdvLoRA is adapted under PGD-3. The experimental results indicate that (1) AdvLoRA consistently outperforms the baseline model, LoRA, across various attack types. This trend continues with the FGSM attack, where AdvLoRA scores 84.44 and 72.21, compared to LoRA’s 61.01 and 54.41. (2) AdvLoRA also demonstrates strong performance under black-box attacks. For the SA, AdvLoRA attains a TR mean score of 87.40 and an IR mean score of 75.83, highlighting its robustness against unseen attack strategies. The results collectively affirm that AdvLoRA not only enhances robustness through adversarial adaptation but also generalizes effectively across different attack types, thereby reinforcing the model’s applicability in real-world scenarios where diverse adversarial threats may be encountered.
Scaling to Larger Backbones. To demonstrate the scalability of the proposed method, we evaluated various sizes of BLIP as the backbone on Flickr30K, as shown in Tab. 9. The findings indicate that: (1) Consistent Performance Improvement. AdvLoRA consistently outperforms the baseline model, LoRA, across both TR and IR metrics, achieving notable scores for different backbone sizes. (2) Enhanced Capability Utilization. This trend continues with the larger BLIP architectures, where AdvLoRA maintains a clear performance edge over LoRA. (3) Significant Gains in Larger Architectures. When scaling to the larger BLIP-2-opt-2.7b backbone, AdvLoRA shows substantial improvements, reaffirming its effectiveness. These results highlight AdvLoRA’s robustness and adaptability, confirming its ability to leverage larger architectures for enhanced retrieval performance. This scalability not only improves model efficiency but also supports its application in complex real-world scenarios, where capturing intricate data patterns is crucial.
Table 9. Performance on Flickr30K when scaling to larger backbone networks. “TR” and “IR” donate text-to-image retrieval and image-to-text retrieval.
Backbone Method TR@Mean IR@Mean Mean
BLIP-base LoRA 80.40 73.24 76.82
AdvLoRA 82.83 74.67 78.75
BLIP-large LoRA 81.73 73.01 77.34
AdvLoRA 84.74 75.56 80.15
BLIP-2-opt-2.7b LoRA 83.32 81.18 82.25
AdvLoRA 87.34 83.66 85.50
Performance on the Natural Data. To evaluate the performance of our proposed method on natural images, we assess AdvLoRA and five baselines on MSCOCO, as shown in Tab. 10. The findings indicate: (1) Trade-off in Performance. While adversarial adaptation enhances robustness against specific attacks, it can degrade performance on natural data, suggesting a trade-off between adversarial resilience and generalization. (2) Superior Metrics of AdvLoRA. AdvLoRA achieves a TR mean score of 85.18 and an IR mean score of 79.92, surpassing all baselines due to its ability to learn semantically invariant feature representations. (3) Versatility Across Retrieval Tasks. Compared to methods like BLIP+FFT and UniAdapter, AdvLoRA excels in TR and IR metrics, demonstrating a retrieval mean score of 79.92, underscoring its reliability. These results affirm AdvLoRA’s effectiveness in balancing robustness and performance in natural image retrieval.
Table 10. Performance on Flickr30K when scaling to larger backbone networks. “TR” and “IR” donate text-to-image retrieval and image-to-text retrieval.
Method TR@Mean IR@Mean R@Mean
BLIP+FFT*+Nat. 74.17 69.11 71.65
BLIP+LoRA*+Nat. 85.29 74.75 80.02
BLIP+LP*+Nat. 86.33 75.38 80.86
UniAdapter*+Nat. 75.29 66.89 71.09
Aurora*+Nat. 84.74 73.54 79.15
AdvLoRA*+Nat. 85.18 74.67 79.92
Case Study. To visually demonstrate the effectiveness of AdvLoRA, we present several test cases from the MSR-VTT dataset, as illustrated in Fig. 5. The results show that: (1) Resilience to Adversarial Attacks. In Video168, AdvLoRA retrieves ”He is playing with ball,” accurately reflecting the scene despite adversarial perturbations, while Aurora shows a noticeable decline in performance. (2) Contextual Understanding. For Video8915, AdvLoRA identifies ”Women preparing to cook a roast,” demonstrating its ability to grasp contextual details, whereas Aurora’s output lacks the same level of accuracy. (3) Complex Scene Interpretation. In Video128, AdvLoRA maintains clarity in retrieving a cartoon character preparing to ride a bicycle, contrasting with Aurora’s irrelevant result, highlighting AdvLoRA’s ability to retain semantic integrity amid adversarial challenges. These examples underscore AdvLoRA’s robustness and contextual awareness in dynamic environments.
Figure 5. Case study of MSR-VTT. We sample and visualize eight frames from the videos. The frames with the devil denote that they are under the adversarial attacks.
@@ 5. Conclusion
In this paper, we aim to alleviate the security risks in the Vision-Language Models (VLMs). First of all, we show the vulnerability of VLMs with various adaptation methods under adversarial attacks via extensive experiments. Besides, as the sizes of VLMs increase, simply applying the conventional adversarial adaptation methods to VLMs easily leads to (1) unpromising adversarial robustness and (2) tremendous parameter costs. To tackle these issues, we propose a novel parameter-efficient adversarial adaptation method called AdvLoRA, which incorporates parameter clustering, parameter alignment, and adaptive parameter updates. Extensive experiments validate the effectiveness and efficiency of AdvLoRA, revealing the intrinsic low-rank property that emerges during the adversarial adaptation process. Our technique, which emphasizes clustering reparameterization and parameter alignment, significantly enhances the adaptation process. This work provides a fresh perspective for researchers focused on security within the broader context of Artificial General Intelligence (AGI). Moving forward, we aim to optimize memory and computational resources during the adaptation process and enhance the adversarial robustness of VLMs, particularly against language-based attacks.
@@ Acknowledgements.
This work is supported by the Natural Science Foundation of China under Grant Nos. 72225011, 72434005 and L242400108.